OpenStudy (anonymous):

what is the smallest sloution to the equation x to the fourth - 34x to the 2 +225=0

OpenStudy (shadowfiend):

So, you can factor this into a pair of quadratics. Basically, you have this: \[x^4 - 34x^2 + 225 = 0\] Imagine that \(u = x^2\). Then you have: \[u^2 - 34u + 225 = 0\] So first you solve for \(u\). Then, you have \(u = x^2\), so \(x = \sqrt{u}\). Does that help?